9/11/16
Summary: Light winds, < 5 m/s, out of the W, partly cloudy.
Actions past 24 hours:
To dos:
Sensor Status:
UPDATE: Gary got Mote #2 running again.
T/RH: ok
P: ok
2D Gill: ok, down at S8
csat u,v: ok
csat ldiag: ok
csat w, tc: ok
EC150: ok
motes: Mote #2 down since yesterday (no soil at #1 and #3)
Wetness: down
radiation: ok, except Mote #2
Tsoil: down
Gsoil: down
Qsoil: down
Cvsoil: not responding
Vbatt: ok, battery at site 15 getting low at night, will monitor